2+2=5 show the working

2+2=5


show the working


 

Grade:9

1 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

Let a & b each be equal to 1. Since a ^ b are equal,

b^2 = ab ...(eq.1)

Since a equals itself, it is obvious that

a^2 = a^2 ...(eq.2)

Subtract equation 1 from equation 2. This yeilds

(a^2) - (b^2) = (a^2)-ab ...(eq. 3)

We can factor both sides of the equation; (a^2)-ab equals a(a-b). Likewise, (a^2)-(b^2) equals (a + b)(a - b) (Nothing fishy is going on here. Ths statement is perfectly true. Plug in numbers and see for yourself!) Substituting into the equation 3 , we get

(a+b)(a-b) = a (a-b) ...(eq.5)

So far, so good. Now divide both sides of the equation by (a-b) and we get

a + b = a ...(eq.5)

b = 0 ...(eq.6)

But we set b to 1 at the very beginning of this proof, so this means that

1 = 0 ...(eq.7)

 

This gives 5=4

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com